OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? 3, 4, 5 28, 26, 12 17, 14, 6 all of the above

OpenStudy (taylorclaudio):

i think it is b

OpenStudy (welshfella):

Check which fit the Pythagoras theorem C^2 = a^2 + b^2 where c = longest side

OpenStudy (welshfella):

Have you tried a?

OpenStudy (anonymous):

3,4, and 5 only, i think - 85% sure of it.

OpenStudy (anonymous):

i am right now but i thought it would be c